OpenStudy (anonymous):

Find the area of a sector that has an arc length of 6 feet on a circle that has a diameter of 4 feet

OpenStudy (anonymous):

\[arc length = \theta/360 *2* \pi*r\] \[area of sector = \theta/360*\pi*r ^{2}\]